The de Broglie wavelength of a tennis ball of mass 60 g moving with a velocity of 10 metres per second is approximately

Grade:12

2 Answers

Sunil Kumar FP
9 years ago
given the mass of the body=60g=60/1000 kg =.06kg
velocity of the ball =10m/s
let the de broglie wavelength be lambda
lambda=h/mv
=6.63*10^-34/(.06*10)
=11.05*10^-34 m
Thus we can see that the debroglie wvelength is very small to meausre.
Damanpreet Singh
6 years ago
When we find wavelength of tennis ballIt to too small And No instrument is available to measure this small amount of wavelength